hello can u tell me about real value of (-2) ^(2/6) .Whether the expression is real or not..... also comment about (-2) ^(1/3)

hello 


can u tell me about real value of (-2) ^(2/6) .Whether the expression is real or not.....


also comment about (-2) ^(1/3) 

Grade:

2 Answers

Chetan Mandayam Nayakar
11 years ago

(-2) ^(2/6)=(-2) ^(1/3)=e(ln(2)+i(2n+1)π)/3, where n is ANY integer

jitendra tiwari
11 years ago

If a^(m/n) is nth root of a^m.............Considering this statement to be true

can we say (-2)^ (2/6) will be 6th root of (-2)^2.....

and can we say that(-2)^(1/3) is 3rd root of -2= negative of third root of 2